OpenStudy (nudeoua):

In the function f(x) = 4(x^2 − 6x +___ ) + 20, what number belongs in the blank to complete the square?

OpenStudy (freckles):

hint: \[x^2+kx+(\frac{k}{2})^2=(x+\frac{k}{2})^2\] you have \[x^2-6x+...\] compare this to the left hand side of the equation I wrote

OpenStudy (freckles):

and whatever you add in you subtract out